# Gaussian Process Tilted Nonparametric Density Estimation using Fisher Divergence Score Matching

### v1

@@ III-A1 Learning base measure parameters $\mu$ and $\Sigma$
For data $x\in\mathbb{R}^{D}$, this matrix inverse is $\mathcal{O}(D^{6})$. This is not prohibitive for the low dimensional problems we focus on. As the dimensionality increases, other algorithms can be used to solve for $\Sigma$. When optimizing Equation 6, the base parameters can be learned by iterating between updates to $\theta$ (Algorithm 2) and $\mu$ and $\Sigma$ (Equations 12 & 16), with slower runtime because the algorithm is now iterative.
@@ III-B Noise Conditional Score Matching
It has often been observed that the approach of Section III-A will have difficulty accurately learning $q_{\theta}$ in regions where little or no data exists, which can lead to inflated density values in low data regions as a learning artifact. To address this, noise conditional score matching adds Gaussian noise of arbitrary variance to the data and learns a noise-dependent score model. A single distribution is learned to approximate the data density in all noise regimes, which provides greater information to the model for learning low density regions.
To this end, the noise-added data model is
[…]
Below, we will write a sample $y\sim p(y|\sigma)$ as $y=x+\xi$ where $x\sim\sum_{i=1}^{N}\frac{1}{N}\delta_{x_{i}}$ and $\xi\sim\mathcal{N}(0,\sigma^{2}I)$. We modify our density to incorporate the parameter $\sigma$ as follows,
@@
$\phi_{\sigma}(y)$ $=$ $\sqrt{2/S}\cos(Zy/\gamma_{\sigma}+c)$ (18)
$\phi^{\prime}_{\sigma}(y)$ $=$ $-\sqrt{2/S}\sin(Zy/\gamma_{\sigma}+c)$ (19)
$\gamma_{\sigma}\coloneqq\sqrt{\gamma^{2}+\sigma^{2}},\qquad\Sigma_{\sigma}=\Sigma+\sigma^{2}I$ (20)
We observe that the noise added to $x$ is naturally incorporated in $q_{\theta}$ in two places: 1) By expanding the variance of the base Gaussian measure to expand the data region, and 2) by increasing the kernel width to learn the Gaussian process at multiple scales. Using the shared $\theta$ learned for all $\sigma>0$, we can then evaluate $q_{\theta}$ at arbitrary resolution, e.g., $q_{\theta}(x)=q_{\theta}(y|\sigma=0)$ since $\phi_{0}(y)=\phi(x)$.
With this modification, the standard noise conditional Fisher divergence objective is
$\begin{array}[]{rl}\mathcal{L}_{\sigma}=&\frac{1}{2}\sum_{i=1}^{N}\mathbb{E}_{p(\xi|\sigma)}\big{[}\|\nabla_{x}\ln q_{\theta}(x_{i}+\xi)\|^{2}\big{]}\\ &+\sum_{i=1}^{N}\mathbb{E}_{p(\xi|\sigma)}\big{[}\mathrm{trace}\big{(}\nabla_{xx}^{2}\ln q_{\theta}(x_{i}+\xi)\big{)}\big{]},\end{array}$ (21)
where the derivatives in our case are given in Equations 7 and 8. We next show that these expectations can be solved analytically and a closed form solution exists for solving $\theta_{\mathrm{FD}\sigma}=\arg\min_{\theta}\mathbb{E}_{p(\sigma)}[\mathcal{L}_{\sigma}]+\frac{\lambda}{2}\theta^{\top}\theta$, for a discretized distribution $\sigma\sim p(\sigma)$ defined later.
@@ III-B1 Derivation of expectations
From Equations 7, 8 and 21, four unique expectations arise. For the expectation of $\cos({z^{\top}y}/{\gamma_{\sigma}}+c)$, where $y=x+\xi$ and $\xi\sim\mathcal{N}(0,\sigma^{2}I)$, [^1] To reduce notation, we write $y$ instead of $y_{\sigma}$. The accompanying $\gamma_{\sigma}$ will indicate the value of $\sigma$ used to generate $y$. we represent the argument as follows,
${z^{\top}y}/{\gamma_{\sigma}}+c~{}~{}\stackrel{{\scriptstyle d}}{{=}}~{}~{}\epsilon+\chi_{\sigma},$
@@
$\chi_{\sigma}\coloneqq{z^{\top}x}/{\gamma_{\sigma}}+c,\quad\epsilon\sim\mathcal{N}(0,{\frac{\sigma^{2}}{\gamma^{2}+\sigma^{2}}}\|z\|^{2}).$
Letting $j=\sqrt{-1}$, the MGF of a univariate Gaussian can be used to derive the following expectation,
$\mathbb{E}[\cos(\epsilon+\chi_{\sigma})]$ $=$ $\mathrm{Re}\left(\mathrm{e}^{j\chi_{\sigma}}\mathbb{E}_{p(\epsilon|\sigma)}\left[\mathrm{e}^{j\epsilon}\right]\right)$
$=$ $\exp\left(-\frac{\|z\|^{2}}{2}\frac{\sigma^{2}}{\gamma^{2}+\sigma^{2}}\right)\cos(\chi_{\sigma}).$
@@
Similarly, $\mathbb{E}[\sin(\epsilon+\chi_{\sigma})]$ is solved using the imaginary component by replacing $\cos$ in Equation III-B1 with $\sin$.
@@
A third expectation arises of the form $\mathbb{E}[\phi_{\sigma s}^{\prime}\phi_{\sigma s^{\prime}}^{\prime}]$ We use the identity $\sin(\alpha)\sin(\beta)=\frac{1}{2}(\cos(\alpha-\beta)-\cos(\alpha+\beta))$, where $\alpha={z_{s}^{\top}\xi}/{\gamma_{\sigma}}+\chi_{\sigma s}$ and $\beta={z_{s^{\prime}}^{\top}\xi}/{\gamma_{\sigma}}+\chi_{\sigma s^{\prime}}$, and again $\chi_{\sigma s}\coloneqq{z_{s}^{\top}x}/{\gamma_{\sigma}}+c_{s}$. The expectation of these cosines follows a similar derivation as above and results in,
$\begin{array}[]{c}\mathbb{E}[\phi^{\prime}_{\sigma s}(y)\phi^{\prime}_{\sigma s^{\prime}}(y)]=\\ \qquad\qquad\quad\frac{1}{S}\exp\Big{(}-\frac{\sigma^{2}\|z_{s}-z_{s^{\prime}}\|^{2}}{2(\gamma^{2}+\sigma^{2})}\Big{)}\cos(\chi_{\sigma s}-\chi_{\sigma s^{\prime}})\\ \qquad\quad~{}~{}~{}~{}~{}~{}-\frac{1}{S}\exp\Big{(}-\frac{\sigma^{2}\|z_{s}+z_{s^{\prime}}\|^{2}}{2(\gamma^{2}+\sigma^{2})}\Big{)}\cos(\chi_{\sigma s}+\chi_{\sigma s^{\prime}})\end{array}$
@@
for all $s,s^{\prime}\in\{1,\dots,S\}$, including $s=s^{\prime}$. Using the identity $\cos(\alpha\pm\beta)=\cos\alpha\cos\beta\mp\sin\alpha\sin\beta$, this can be simplified for faster computation as shown later.
@@
A final expectation is of the cross term in the square of the noise-added Equation 7. This expectation takes the form $\mathbb{E}_{p(\xi|\sigma)}[\xi^{\top}\Sigma_{\sigma}^{-1}z\sin(\xi^{\top}z/\gamma_{\sigma}+\chi_{\sigma})]$. We solve this by writing vector products in sum form over the dimensions of $\xi$ and solving $D$ separate expectations. Using the shorthand $g(\xi)={\xi^{\top}z}/{\gamma_{\sigma}}+\chi_{\sigma}$, and recalling that $\chi_{\sigma}$ is also a function of $z$, we solve by using the equality
$\mathbb{E}\left[\xi_{d}\sin g(\xi)\right]=-x_{d}\mathbb{E}\sin g(\xi)-\gamma_{\sigma}({\partial}/{\partial z_{d}})\mathbb{E}\cos g(\xi).$
@@
After cancellations, and using the previous expectations for $\sin$ and $\cos$, the total expectation is
$\textstyle\sum_{d=1}^{D}(\Sigma_{\sigma}^{-1}z)_{d}\mathbb{E}\left[\xi_{d}\sin(\xi^{\top}z/\gamma_{\sigma}+\chi_{\sigma})\right]=$ (23)
$\textstyle\frac{\sigma^{2}}{\gamma_{\sigma}}(z^{\top}\Sigma_{\sigma}^{-1}z)\exp\Big{(}-\frac{\|z\|^{2}}{2}\frac{\sigma^{2}}{\gamma^{2}+\sigma^{2}}\Big{)}\cos\Big{(}\frac{z^{\top}x}{\gamma_{\sigma}}+c\Big{)}.$
@@
We summarize all relevant expectations in vector form in Table I, where we use the following definitions:
$\delta_{\sigma}(s)$ $=$ $\exp(-\textstyle\frac{1}{2}(\sigma/\gamma_{\sigma})^{2}\|z_{s}\|^{2}),$ (24)
$\Delta^{+}_{\sigma}(s,s^{\prime})$ $=$ $\exp(-\textstyle\frac{1}{2}(\sigma/\gamma_{\sigma})^{2}\|z_{s}+z_{s^{\prime}}\|^{2}),$ (25)
$\Delta^{-}_{\sigma}(s,s^{\prime})$ $=$ $\exp(-\textstyle\frac{1}{2}(\sigma/\gamma_{\sigma})^{2}\|z_{s}-z_{s^{\prime}}\|^{2}),$ (26)
@@
where $\delta_{\sigma}$ is a vector and $\Delta_{\sigma}^{+}$ and $\Delta_{\sigma}^{-}$ are matrices. Again, our definitions described previously are $\gamma_{\sigma}=\sqrt{\gamma^{2}+\sigma^{2}}$, $\|ZM\|^{2}\equiv(\|Mz_{1}\|^{2},\dots,\|Mz_{S}\|^{2})^{\top}$ and $\Sigma_{\sigma}=\Sigma+\sigma^{2}I$. As previously defined in Equation 18, the RFF vector $\phi_{\sigma}$ indicates that the kernel width $\gamma_{\sigma}$ is used, while $(z,c)$ are shared for all values of $\sigma$.
TABLE I: List of Expectations (also see Eqs. 18,19,2426)
$\mathbb{E}[\phi_{\sigma}(y)]$ $=$ $\delta_{\sigma}\odot\phi_{\sigma}(x)$
$\mathbb{E}[\phi^{\prime}_{\sigma}(y)]$ $=$ $\delta_{\sigma}\odot\phi^{\prime}_{\sigma}(x)$
$\mathbb{E}[\phi^{\prime}_{\sigma}(y)\phi^{\prime}_{\sigma}(y)^{\top}]$ $=$ $\frac{1}{2}(\Delta^{-}_{\sigma}-\Delta^{+}_{\sigma})\odot\phi_{\sigma}(x)\phi_{\sigma}(x)^{\top}$
$+$ $\frac{1}{2}(\Delta^{-}_{\sigma}+\Delta^{+}_{\sigma})\odot\phi^{\prime}_{\sigma}(x)\phi^{\prime}_{\sigma}(x)^{\top}$
$\mathbb{E}[(Z\Sigma_{\sigma}^{-1}\xi)\odot\phi^{\prime}_{\sigma}(y)]$ $=$ $-\frac{\sigma^{2}}{\gamma_{\sigma}}\delta_{\sigma}\odot\|Z\Sigma_{\sigma}^{-\frac{1}{2}}\|^{2}\odot\phi_{\sigma}(x)$
III-B2 Algorithm summary and discussion
We can now calculate the noise conditional objective function to be optimized. We use a uniform continuous distribution on $\sigma\in[\frac{\gamma}{D},\sigma_{\mathrm{max}}+\frac{\gamma}{D}]$ approximated at $H$ evenly spaced points starting at $\gamma$ and terminating at $\frac{H-1}{H}\sigma_{\mathrm{max}}+\frac{\gamma}{D}$. The value of $\sigma_{\mathrm{max}}$ is a parameter to be set. The minimum value $\frac{\gamma}{D}\ll\sigma_{\mathrm{max}}$ is the kernel width, which we empirically found resolves GP fluctuation issues that occasionally occur in large regions of no data when $\sigma\rightarrow 0$. Without the prior,
$\mathcal{L}~{}=~{}\mathbb{E}_{p(\sigma)}[\mathcal{L}_{\sigma}]~{}\approx~{}\frac{1}{H}\sum_{h=1}^{H}\mathcal{L}_{\frac{\gamma}{D}+{\frac{h-1}{H}\sigma_{\mathrm{max}}}}.$ (27)
The gradient $\mathcal{L}_{\sigma}$ for each discretized value of $\sigma$ is
$\nabla_{\theta}\mathcal{L}_{\sigma}=\textstyle-\frac{1}{\gamma_{\sigma}^{2}}\|Z\|^{2}\odot\sum_{i=1}^{N}\mathbb{E}[\phi_{\sigma}(x_{i}+\xi_{i})]$ (28)
$\textstyle\qquad\quad~{}-\frac{1}{\gamma_{\sigma}}\sum_{i=1}^{N}\mathbb{E}[\phi_{\sigma}^{\prime}(x_{i}+\xi_{i})]\odot Z\Sigma_{\sigma}^{-1}(x_{i}-\mu)$
$\textstyle\qquad\quad~{}-\frac{1}{\gamma_{\sigma}}\sum_{i=1}^{N}\mathbb{E}\left[(Z\Sigma_{\sigma}^{-1}\xi_{i})\odot\phi_{\sigma}^{\prime}(x_{i}+\xi_{i})\right]$
$+\textstyle\Big{(}\frac{1}{\gamma_{\sigma}^{2}}ZZ^{\top}\odot\,\sum_{i=1}^{N}\mathbb{E}[\phi_{\sigma}^{\prime}(x_{i}+\xi_{i})\phi_{\sigma}^{\prime}(x_{i}+\xi_{i})^{\top}]\Big{)}\theta.$
These terms parallel Equation 10 for the noise-free case. Using the expectations in Table I and including the Gaussian prior regularization of $\theta$ with parameter $\lambda$, the closed-form solution for $\theta$ that minimizes $\frac{1}{H}\sum_{h=1}^{H}\mathcal{L}_{\frac{\gamma}{D}+{\frac{h-1}{H}\sigma_{\mathrm{max}}}}+\frac{\lambda}{2}\theta^{\top}\theta$ is given in Algorithm 3. We observe that Algorithm 2 results as a special case when $\sigma\in\{0\}$.
@@ III-B2 Algorithm summary and discussion
6:Def. $\|ZM\|^{2}=(\|Mz_{1}\|^{2},\dots,\|Mz_{S}\|^{2})^{\top}$ $M\in\mathbb{S}_{++}^{d}$
7:Define $\gamma_{\sigma}=\sqrt{\gamma^{2}+\sigma^{2}}$ and $\Sigma_{\sigma}=\Sigma+\sigma^{2}I$ for values of $\sigma$ in the set $\Omega_{\sigma}=\left\{\frac{\gamma}{D}+\frac{h-1}{H}\sigma_{\mathrm{max}}:h=1,\dots,H\right\}$
8:For data $\{x_{i}\}_{i=1:N}$ and each $\sigma\in\Omega_{\sigma}$, calculate RFF vectors $\phi_{\sigma}(x_{i})$ and $\phi^{\prime}_{\sigma}(x_{i})$ as in Eqs 18 & 19, and
@@
III-C Variational Predictive Distribution
As an alternative approach, we propose modeling uncertainty in $\theta$ and constructing the predictive distribution $p(x|x_{1},\dots,x_{N})$, with the same goal of eliminating artificially large densities in $p_{\theta}(x)$ that result from a point estimate of $\theta$. Since the log-linear form of our score uniquely allows for the predictive distribution to be approximated in closed form, it provides an alternative to the noise conditional framework used to learn deeper score functions.
To this end, we observe that the quadratic form in $\theta$ of the Fisher divergence in Equation 6, combined with the Gaussian prior on $\theta$ suggests a Gaussian posterior for $\theta$. Since the Fisher divergence is not a proper log likelihood, we instead approximate the posterior by maximizing a variational ELBO-type function,
$\mathcal{L}(q(\theta))=\mathbb{E}_{q(\theta)}\left[-\frac{1}{2\eta}\mathbb{E}_{p(x)}\Big{\|}\nabla_{x}\ln\frac{p(x)}{q_{\theta}(x)}\Big{\|}^{2}+\ln\frac{p(\theta)}{q(\theta)}\right].$
Here, the log likelihood of the typical ELBO is replaced with the negative Fisher divergence. We also include an additional variational tempering (VT) parameter $\eta$, originally proposed by Mandt et al. [32] to anneal the ELBO for better local optimal learning of $q(\theta)$. In VT, $\eta$ decreases to $1$ as a function of iteration to ensure optimization of the desired ELBO; since we will be able to solve for the optimal $q(\theta)$ in closed form, we instead use $\eta$ to perform a necessary regularization to balance the Fisher and KL divergence terms, which we discuss later.
Following the generic derivations of variational inference theory [33], the optimal $q(\theta)$ is

### v2

@@ III-A1 Learning $\mu$ and $\Sigma$ of base measure
For data $x\in\mathbb{R}^{d}$, this matrix inverse is $\mathcal{O}(d^{6})$. This is not prohibitive for the low dimensional problems we consider. As the dimensionality increases, other algorithms can be used to solve for $\Sigma$. When optimizing Equation 6, the base parameters can be learned by iterating between updates to $\theta$ (Algorithm 2) and $\mu$ and $\Sigma$ (Equations 12 & 16). Because the algorithm is now iterative, it takes significantly more time, which is a major drawback. We set $\mu$ and $\Sigma$ as previously described.
@@ III-B Noise Conditional Score Matching
It has been observed that the approach of Section III-A can have difficulty accurately learning $q_{\theta}$ in regions where little or no data exists. This can lead to inflated density values in low data regions as a learning artifact, especially for GP-based scores, since reversion to the model prior in these regions produce random functions with no preference between high or low density. To address this learning issue, noise conditional score matching adds Gaussian noise of arbitrary variance to the data and learns a noise-dependent score model. A single distribution is learned to approximate the data density in all noise regimes, which provides greater information to the model for learning low density regions.
The noise-added model modifies the data distribution as
$p(y|\sigma)=\int p(y|x,\sigma)p(x)dx\approx\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(y|x_{i},\sigma^{2}I).$
Below, we will equivalently write a sample $y\sim p(y|\sigma)$ as $y=x+\xi$ where $x\sim\sum_{i=1}^{N}\frac{1}{N}\delta_{x_{i}}$ and $\xi\sim\mathcal{N}(0,\sigma^{2}I)$.
Next, we need to modify our TGP density to account for the noise level $\sigma$. In other score models this is simply done by including $\sigma$ as an input to the neural network. The Gaussian process permits a more natural probabilistic adaptation,
@@
$\phi_{\sigma}(y)\hskip-2.0pt$ $=$ $\hskip-2.0pt\sqrt{2/S}\cos(Zy/\gamma_{\sigma}+c),$ (18)
$\gamma_{\sigma}\coloneqq\sqrt{\gamma^{2}+\sigma^{2}},\quad\Sigma_{\sigma}=\Sigma+\sigma^{2}I.$ (19)
We observe that the noise added to $x$ is naturally incorporated in $q$ in two places: 1) By expanding the variance of the base Gaussian measure to expand the data region of interest, and 2) by increasing the kernel width to learn a multiscale GP. Using one $\theta$ learned for all $\sigma\geq 0$, we can evaluate $q_{\theta}$ at arbitrary resolution, including the original $q_{\theta}(x)$ when $\sigma=0$.
With this modification, the noise conditional Fisher divergence using parameter $\sigma$ is
$\hskip-12.0pt\begin{array}[]{rl}\mathcal{L}_{\sigma}=&\displaystyle\tfrac{1}{2}\sum\nolimits_{i=1}^{N}\mathbb{E}_{p(\xi|\sigma)}\big[\|\nabla_{x}\ln q_{\theta}(x_{i}+\xi)\|^{2}\big]\\ &\displaystyle+\sum\nolimits_{i=1}^{N}\mathbb{E}_{p(\xi|\sigma)}\big[\mathrm{trace}\big(\nabla_{xx}^{2}\ln q_{\theta}(x_{i}+\xi)\big)\big].\end{array}\hskip-10.0pt$ (20)
The form of the derivatives were given in Equations 7 and 8, except the new $\sigma$-dependent $\phi_{\sigma}=-\phi_{\sigma}^{\prime\prime}$ of Equation 18 is used instead, along with its derivative
$\phi^{\prime}_{\sigma}(y)=-\sqrt{2/S}\sin(Zy/\gamma_{\sigma}+c).$ (21)
We next show that these expectations can be solved analytically and a closed form solution exists for solving the total objective $\theta_{\mathrm{FD}\sigma}=\arg\min_{\theta}\mathbb{E}_{p(\sigma)}[\mathcal{L}_{\sigma}]+\frac{\lambda}{2}\theta^{\top}\theta$ for some preset discretized distribution $\sigma\sim p(\sigma)$ defined later.
@@ III-B1 Derivation of expectations
From Equation 20 four unique expectations arise. For the expectation of $\cos({z^{\top}y}/{\gamma_{\sigma}}+c)$, where $y=x+\xi$ and $\xi\sim\mathcal{N}(0,\sigma^{2}I)$, we represent the input to the cosine as follows,
${z^{\top}(x+\xi)}/{\gamma_{\sigma}}+c~~\stackrel{{\scriptstyle d}}{{=}}~~\epsilon+\chi_{\sigma},$
@@
$\chi_{\sigma}\coloneqq{z^{\top}x}/{\gamma_{\sigma}}+c,\quad\epsilon\sim\mathcal{N}\left(0,{\tfrac{\sigma^{2}}{\gamma^{2}+\sigma^{2}}}\|z\|^{2}\right).$
The MGF of a univariate Gaussian can then be used to derive the following expectation,
$\mathbb{E}[\cos(\epsilon+\chi_{\sigma})]\hskip-5.0pt$ $=$ $\hskip-5.0pt\mathrm{Re}\left(\mathrm{e}^{j\chi_{\sigma}}\mathbb{E}_{p(\epsilon|\sigma)}\left[\mathrm{e}^{j\epsilon}\right]\right)$
$=$ $\hskip-5.0pt\exp\left(-\frac{\|z\|^{2}}{2}\frac{\sigma^{2}}{\gamma^{2}+\sigma^{2}}\right)\cos(\chi_{\sigma}),$
@@
where $j=\sqrt{-1}$. Similarly, $\mathbb{E}[\sin(\epsilon+\chi_{\sigma})]$ is solved using the imaginary component, which replaces $\cos$ with $\sin$ above.
@@
A third expectation arises of the form $\mathbb{E}[\phi_{\sigma s}^{\prime}\phi_{\sigma s^{\prime}}^{\prime}]$ We use the identity $\sin(\alpha)\sin(\beta)=\frac{1}{2}(\cos(\alpha-\beta)-\cos(\alpha+\beta))$, where $\alpha={z_{s}^{\top}\xi}/{\gamma_{\sigma}}+\chi_{\sigma s}$ and $\beta={z_{s^{\prime}}^{\top}\xi}/{\gamma_{\sigma}}+\chi_{\sigma s^{\prime}}$ here, and again let $\chi_{\sigma s}\coloneqq{z_{s}^{\top}x}/{\gamma_{\sigma}}+c_{s}$. The expectation of these cosines follows a similar derivation as above and results in
@@
$\begin{array}[]{c}\mathbb{E}[\phi^{\prime}_{\sigma s}(y)\phi^{\prime}_{\sigma s^{\prime}}(y)]=~~~\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\\ \qquad\qquad\quad~~\frac{1}{S}\exp\Big(-\frac{\sigma^{2}\|z_{s}-z_{s^{\prime}}\|^{2}}{2(\gamma^{2}+\sigma^{2})}\Big)\cos(\chi_{\sigma s}-\chi_{\sigma s^{\prime}})\\ \qquad\quad~~~~~~-\frac{1}{S}\exp\Big(-\frac{\sigma^{2}\|z_{s}+z_{s^{\prime}}\|^{2}}{2(\gamma^{2}+\sigma^{2})}\Big)\cos(\chi_{\sigma s}+\chi_{\sigma s^{\prime}})\vskip 6.0pt\end{array}$ for all $s,s^{\prime}\in\{1,\dots,S\}$, including $s=s^{\prime}$. Using the identity $\cos(\alpha\pm\beta)=\cos\alpha\cos\beta\mp\sin\alpha\sin\beta$, this can be simplified for faster computation as shown later.
TABLE I: List of Expectations (see Eqs. 18, 21, 2426)
$\mathbb{E}[\phi_{\sigma}(y)]$ == $\delta_{\sigma}\odot\phi_{\sigma}(x)$
$\mathbb{E}[\phi^{\prime}_{\sigma}(y)]$ == $\delta_{\sigma}\odot\phi^{\prime}_{\sigma}(x)$
$\mathbb{E}[\phi^{\prime}_{\sigma}(y)\phi^{\prime}_{\sigma}(y)^{\top}]$ == $\frac{1}{2}(\Delta^{-}_{\sigma}-\Delta^{+}_{\sigma})\odot\phi_{\sigma}(x)\phi_{\sigma}(x)^{\top}$
++ $\frac{1}{2}(\Delta^{-}_{\sigma}+\Delta^{+}_{\sigma})\odot\phi^{\prime}_{\sigma}(x)\phi^{\prime}_{\sigma}(x)^{\top}$
$\mathbb{E}[(Z\Sigma_{\sigma}^{-1}\xi)\odot\phi^{\prime}_{\sigma}(y)]$ == $-\frac{\sigma^{2}}{\gamma_{\sigma}}\delta_{\sigma}\odot\|Z\Sigma_{\sigma}^{-\frac{1}{2}}\|^{2}\odot\phi_{\sigma}(x)$
@@
A final expectation occurs in the cross term of the square of the gradient in Equation 7. This expectation takes the form $\mathbb{E}_{p(\xi|\sigma)}[\xi^{\top}\Sigma_{\sigma}^{-1}z\sin(\xi^{\top}z/\gamma_{\sigma}+\chi_{\sigma})]$. We solve this by writing vector products in sum form over the $d$ dimensions of $\xi$ and solving separate expectations then adding. Using the shorthand $g(\xi)={\xi^{\top}z}/{\gamma_{\sigma}}+\chi_{\sigma}$, and recalling that $\chi_{\sigma}$ is also a function of $z$, we solve by using the equality
$\mathbb{E}\left[\xi_{l}\sin g(\xi)\right]=-x_{l}\mathbb{E}\sin g(\xi)-\gamma_{\sigma}({\partial}/{\partial z_{l}})\mathbb{E}\cos g(\xi).$
@@
Using the expectations for $\sin$ and $\cos$ from Equation III-B1 and making some cancellations, the total expectation is
$\hskip-23.0pt\sum\nolimits_{l=1}^{d}(\Sigma_{\sigma}^{-1}z)_{l}\mathbb{E}\left[\xi_{l}\sin(\xi^{\top}z/\gamma_{\sigma}+\chi_{\sigma})\right]=$ (23)
$\frac{\sigma^{2}}{\gamma_{\sigma}}(z^{\top}\Sigma_{\sigma}^{-1}z)\exp\Big(-\frac{\|z\|^{2}}{2}\frac{\sigma^{2}}{\gamma^{2}+\sigma^{2}}\Big)\cos\Big(\frac{z^{\top}x}{\gamma_{\sigma}}+c\Big).$
@@
We summarize all expectations in vector form in Table I, where we use the following definitions:
$\delta_{\sigma}(s)\hskip-5.0pt$ $=$ $\hskip-5.0pt\exp(-\textstyle\frac{1}{2}(\sigma/\gamma_{\sigma})^{2}\|z_{s}\|^{2}),$ (24)
$\Delta^{+}_{\sigma}(s,s^{\prime})\hskip-5.0pt$ $=$ $\hskip-5.0pt\exp(-\textstyle\frac{1}{2}(\sigma/\gamma_{\sigma})^{2}\|z_{s}+z_{s^{\prime}}\|^{2}),$ (25)
$\Delta^{-}_{\sigma}(s,s^{\prime})\hskip-5.0pt$ $=$ $\hskip-5.0pt\exp(-\textstyle\frac{1}{2}(\sigma/\gamma_{\sigma})^{2}\|z_{s}-z_{s^{\prime}}\|^{2}),$ (26)
@@
where $\delta_{\sigma}$ is an $S$-dimensional vector and $\Delta_{\sigma}^{+}$ and $\Delta_{\sigma}^{-}$ are $S\times S$ matrices. Again, we use definitions described previously: $\gamma_{\sigma}=\sqrt{\gamma^{2}+\sigma^{2}}$, $\|ZM\|^{2}\equiv(\|Mz_{1}\|^{2},\dots,\|Mz_{S}\|^{2})^{\top}$ and $\Sigma_{\sigma}=\Sigma+\sigma^{2}I$. Also, as defined in Equation 18, the RFF vector $\phi_{\sigma}$ indicates that the kernel width $\gamma_{\sigma}$ is used, while $(z,c)$ are shared for all values of $\sigma$.
@@
6:Define $\|ZM\|^{2}=(\|Mz_{1}\|^{2},\dots,\|Mz_{S}\|^{2})^{\top}$, $M\in\mathbb{S}_{++}^{d}$
7:Define $\gamma_{\sigma}=\sqrt{\gamma^{2}+\sigma^{2}}$ and $\Sigma_{\sigma}=\Sigma+\sigma^{2}I$ for values of $\sigma$ in the set $\Omega_{\sigma}=\left\{\frac{h-1}{H}\sigma_{\mathrm{max}}:h=1,\dots,H\right\}$
8:For data $\{x_{i}\}_{i=1:N}$ and each $\sigma\in\Omega_{\sigma}$, calculate RFF vectors $\phi_{\sigma}(x_{i})$ and $\phi^{\prime}_{\sigma}(x_{i})$ as in Eqs 18 & 21, and
@@
III-B2 Algorithm summary and discussion
We can now calculate the complete noise conditional objective function. We use a uniform distribution on $\sigma\in[0,\sigma_{\mathrm{max}}]$ discretized at $H$ evenly spaced points starting at 0 and terminating at $\frac{H-1}{H}\sigma_{\mathrm{max}}$. The value of $\sigma_{\mathrm{max}}$ is a parameter to be set. Without the prior,
$\mathcal{L}{}~=~\mathbb{E}_{p(\sigma)}[\mathcal{L}_{\sigma}]~=~\tfrac{1}{H}\sum\nolimits_{h=1}^{H}\mathcal{L}_{{\frac{h-1}{H}\sigma_{\mathrm{max}}}}.$ (27)
The gradient of $\mathcal{L}_{\sigma}$ for each value of $\sigma$ is
$\hskip-20.0pt\nabla_{\theta}\mathcal{L}_{\sigma}=-\tfrac{1}{\gamma_{\sigma}^{2}}\|Z\|^{2}\odot\sum\nolimits_{i=1}^{N}\mathbb{E}[\phi_{\sigma}(x_{i}+\xi_{i})]$ (28)
$\quad~\,\,\,-\tfrac{1}{\gamma_{\sigma}}\sum\nolimits_{i=1}^{N}\mathbb{E}[\phi_{\sigma}^{\prime}(x_{i}+\xi_{i})]\odot Z\Sigma_{\sigma}^{-1}(x_{i}-\mu)$
$\quad~\,\,\,-\tfrac{1}{\gamma_{\sigma}}\sum\nolimits_{i=1}^{N}\mathbb{E}\left[(Z\Sigma_{\sigma}^{-1}\xi_{i})\odot\phi_{\sigma}^{\prime}(x_{i}+\xi_{i})\right]$
$~~+\Big(\tfrac{1}{\gamma_{\sigma}^{2}}ZZ^{\top}\odot\,\sum\nolimits_{i=1}^{N}\mathbb{E}[\phi_{\sigma}^{\prime}(x_{i}+\xi_{i})\phi_{\sigma}^{\prime}(x_{i}+\xi_{i})^{\top}]\Big)\theta.$
These terms parallel Equation 10 for the noise-free case. Using the expectations in Table I and including the Gaussian prior regularization of $\theta$ with parameter $\lambda$, the closed form solution for $\theta$ that minimizes $\mathcal{L}+\frac{\lambda}{2}\theta^{\top}\theta$ is given in Algorithm 3. We observe that Algorithm 2 is the special case when $\sigma\in\{0\}$.
III-C Fisher Variational Predictive Distribution
Noise conditional score models overcome uncertainty in low data regions by adding noise to the data. As an alternative, we propose a Bayesian formulation that models uncertainty in $\theta$ and constructs the predictive distribution $p(x|x_{1},\dots,x_{N})$, with the same goal of eliminating high density artifacts in low information regions. Since the log-linear form of the TGP score uniquely allows for a predictive distribution to be approximated in closed form, it provides an appealing alternative here, though noise conditional models remain much easier to work with for deeper score functions.
First, we observe that the fact that Fisher divergence is quadratic in $\theta$, combined with the Gaussian prior on $\theta$, suggests approximating the posterior of $\theta$ with a Gaussian. Since the Fisher divergence does not correspond to a log likelihood, where $\ln q_{\theta}(x)$ would be expected instead, using it for variational inference stretches the Bayesian framework. Nevertheless, we note a tractable ELBO-type function suggests itself for optimization,
$\mathcal{L}(q(\theta))=\mathbb{E}_{q(\theta)}\hskip-3.0pt\left[-\frac{1}{2\eta}\mathbb{E}_{p(x)}\Big\|\nabla_{x}\ln\frac{p(x)}{q_{\theta}(x)}\Big\|^{2}+\ln\frac{p(\theta)}{q(\theta)}\right].$
We see the log likelihood of the typical ELBO is replaced with the negative Fisher divergence. We also include an additional variational tempering (VT) parameter $\eta$, originally proposed by Mandt et al. [mandt2016variational] to anneal the ELBO for better local optimal learning of $q(\theta)$. In VT, $\eta$ decreases to $1$ as a function of iteration to ensure optimization of the desired ELBO; since we will be able to find the optimal $q(\theta)$ in closed form, we instead use $\eta$ to perform a necessary regularization to balance the Fisher and KL divergence terms in $\mathcal{L}$.
Following the VI theory [bishop2006pattern], the $q(\theta)$ that maximizes $\mathcal{L}$ is